OpenStudy (anonymous):

A circular sundial has a circumference of 32􏰈 inches. What is the area of the sundial?

OpenStudy (photon336):

\[A = pir ^{2}\] we know that our circumference is 32 inches so we must find the radius to find the area. \[C = 2pir\] \[\frac{ c }{ 2\pi } = r \] \[\frac{ 32 inches }{ 2\pi } = radius \]

OpenStudy (anonymous):

Edit: A circular sundial has a circumference of 32π􏰈 inches. What is the area of the sundial?

OpenStudy (photon336):

ok @frebrezze let me modify my answer

OpenStudy (photon336):

\[\frac{ 32\pi }{ 2\pi } = 16 inches = r\]

OpenStudy (photon336):

now \[a = pir ^{2}\] \[\pi*(16)^{2} = 256\pi \]

OpenStudy (photon336):

does this make sense? @frebrezze ?

OpenStudy (anonymous):

Thank you!!! sorry for typing it wrong and yes

OpenStudy (photon336):

No problem and while we're at it Welcome to Open Study @frebrezze